OpenStudy (anonymous):

3/4b is greater than or equal to -9/8

OpenStudy (anonymous):

3/4 ≥ -9/8

OpenStudy (anonymous):

Multiply both sides by 4/3 the right side looks like: -9/8 * 4/3 = -36/24 = -3/2 so b ≤ -3/2

OpenStudy (anonymous):

looks a typo above ^^^ it should be: \[b \ge -\frac{3}{2}\]

OpenStudy (anonymous):

oops b ≥ -3/2 sorry typo